OpenStudy (anonymous):

help!!!! differentiate lnx+3x^2

OpenStudy (anonymous):

1/x + 6x

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

1/x + 6x

OpenStudy (amistre64):

Dx [ln(x) + 3x^2] = Dx [ln(x)] + Dx [3x^2] = Dx [ln(x)] + 6x = 1/x + 6x

jimthompson5910 (jim_thompson5910):

Is it \[\large \ln(x)+3x^2\] or is it \[\large \ln(x+3x^2)\]

OpenStudy (anonymous):

first one !!!

jimthompson5910 (jim_thompson5910):

then the derivative is \[\large \frac{1}{x}+6x\], which means you have the correct answer.

OpenStudy (anonymous):

thnx guys !!!